OpenStudy (anonymous):

solve by graphing 2x=y+1 2x-y=5

OpenStudy (anonymous):

they wont have any solutions lfthese are straigth line equation and parallel lines

OpenStudy (anonymous):

how do the lines loook like

OpenStudy (anonymous):

jackie they shld be shown on a graph .. would be better to explain u

OpenStudy (anonymous):

there not shown on a graph they just give that equation and say to graph it

OpenStudy (anonymous):

where would the points go

OpenStudy (anonymous):

look these are two lines which should be drawn tats y i was askin u to come there

OpenStudy (anonymous):

i was there u werent

OpenStudy (anonymous):

come back i was on some other question

OpenStudy (anonymous):

send me the link again

OpenStudy (anonymous):

solve the equations into "y = mx + b" format. b is the y-intercept. m is the slope. Using this info, you should be able to draw the 2 lines on a graph, and the solution is where they cross.

OpenStudy (anonymous):

the lines are parallel because they have same gradient

OpenStudy (anonymous):

My bad. Nick is right. When you solve for y, you get y = 2x - 1 and y = 2x - 5. You can graph those, and you'll see they're parallel because they have the same slope. So there is no solution. You can also see this by substituting "y+1" for "2x" in the second equation, which makes "y + 1 - y = 5," which simplifies to "1 = 5," which is impossible.